# Rise and Fall of Anderson Localization By Lattice Vibrations: A Time-Dependent Machine Learning Approach

### v1

@@ Appendix B Clustering in the Frozen Approximation
Figure B.1: Phase diagram of LSCO in the frozen approximation. The phase diagram was derived using machine learning-based clustering to analyze time series data of the wavefunction evolution within systematically varied time-independent deformation potentials.
@@
Figure B.1 shows the phase diagram for LSCO derived from clustering in the frozen approximation. As in the dynamic case, three phases are identified. Supported by the arguments presented in the main text and especially Fig. 6, Phases I and II are very similar to the dynamic case as the electron dynamics are highly adiabatic in these regimes. The key difference emerges in Phase III, the Anderson localization phase. Here, strong localization due to quantum interference in the static potential field leads to complete Anderson localization, contrasting sharply with the transient localization observed in the dynamic case.

### v2

@@ Appendix B Clustering in the Frozen Approximation
Figure B.1 shows the phase diagram for LSCO derived from clustering in the frozen approximation. As in the dynamic case, three phases are identified. Supported by the arguments presented in the main text and especially Figure 6, Phases I and II are very similar to the dynamic case as the electron dynamics are highly adiabatic in these regimes. The key difference emerges in Phase III, the Anderson localization phase. Here, strong localization due to quantum interference in the static potential field leads to complete Anderson localization, contrasting sharply with the transient localization observed in the dynamic case.
Figure B.1: Phase diagram of LSCO in the frozen potential field. The phase diagram was derived using a machine learning-based clustering algorithm to analyze time series data of the wavefunction evolution within systematically varied deformation potentials. This analysis involved variations in temperature $\tilde{T}$ and effective coupling $\tilde{G}$; every point corresponds to one unique configuration. The following three clusters were identified: (I) refractive scattering region (green), (II) diffraction behavior (blue), and (III) Anderson localization (red). The size of the points indicates the level of agreement across an ensemble of different wavefunction data sets.